Hello
17x+9 = 12x² + 14x-6
How would I simplify this to discover X's value
Thank you

Answer by ankor@dixie-net.com(12692)

17x + 9 = 12x² + 14x - 6
;
Group everything on the left
0 = 12x^2 + 14x - 17x - 6 - 9
:
A quadratic equation:
12x^2 - 3x - 15 = 0
:
Simplify, divide by 3:
4x^2 - x - 5 = 0
Factors to:
(4x-5)(x+1) = 0
:
4x = 5
x = 5/4
x = 1.25
and
x = -1
:
:
Check solution in original equation using x = 1.25
17(1.25) + 9 = 12(1.25^2) + 14(1.25) - 6
21.25 + 9 = 18.75 + 17.5 - 6
30.25 = 30.25
:
You can check the x=-1 solution in the same way